# DogLegs: Robust Proprioceptive State Estimation for Legged Robots Using Multiple Leg-Mounted IMUs

### v1

@@ Abstract
Robust and accurate proprioceptive state estimation of the main body is crucial for legged robots to execute tasks in extreme environments where exteroceptive sensors, such as LiDARs and cameras may become unreliable. In this paper, we propose DogLegs, a state estimation system for legged robots that fuses the measurements from a body-mounted inertial measurement unit (Body-IMU), joint encoders, and multiple leg-mounted IMUs (Leg-IMU) using an extended Kalman filter (EKF). The filter system contains the error states of all IMU frames. The Leg-IMUs are used to detect foot contact, thereby providing zero velocity measurements to update the state of the Leg-IMU frames. Additionally, we compute the relative position constraints between the Body-IMU and Leg-IMUs by the leg kinematics and use them to update the main body state and reduce the error drift of the individual IMU frames. Field experimental results have shown that our proposed system can achieve better state estimation accuracy compared to the traditional leg odometry method (using only Body-IMU and joint encoders) across different terrains. We make our datasets publicly available to benefit the research community.
@@ I Introduction
Current approaches to this problem primarily use the robot body-mounted IMU (Body-IMU), joint encoders and foot force sensors [1, 8, 4, 9]. In these systems, the Body-IMU initially propagates the robot’s main body state, which is then refined by integrating leg kinematics estimated from joint encoders upon detecting foot contact using the foot force sensor. The most widely used approach for foot contact detection relies on empirically thresholding force sensor readings, but this method is susceptible to errors due to slippage and sensor degradation over time. In addition, the foot positions during swing phase cannot be estimated [1] and need to be relocated when the foot regains contact with the ground, introducing further error into the robot state estimation.
@@
The main contribution of this paper is the development of DogLegs, a framework that fuses the full state estimation information from all the individual IMU systems, including the Body-IMU and multiple Leg-IMUs, through an EKF. Specifically, each IMU system independently performs state propagation while a comprehensive error-state vector is maintained to account for errors across all IMU frames. The Leg-IMUs are used for foot contact detection, providing zero velocity measurements to update the state of the Leg-IMU frames. Additionally, we use leg kinematics to compute the relative position constraints between the Body-IMU and Leg-IMUs. This measurement helps update the main body state, reduce error drift in individual IMU frames, and estimate and compensate for IMU bias errors. Fig. 1 illustrates the installation of multiple Leg-IMUs and the coordinate systems involved within our proposed DogLegs framework, using a quadruped robot as an example.
@@ II-A Proprioceptive Legged Robot State Estimation
An early approach to proprioceptive state estimation for legged robots was proposed by Bloesch et al. [1], who employed an EKF to fuse outputs from a Body-IMU, joint encoders, and foot force sensors. This method augmented the foot contact position into the state and updated it with the main body state using the leg kinematics when the foot contacts ground. This approach has become the basis for many subsequent state estimation frameworks. Later, Bloesch et al. [8] extended this work by incorporating the zero velocity measurement and replacing the EKF with an unscented Kalman filter to improve accuracy and robustness.
@@ III Our Approach
Fig. 2 illustrates the algorithmic structure of our proposed DogLegs system. Our method is implemented using an error-state EKF to fuse the measurements from joint encoders and all the IMUs. The Leg-IMUs propagate the states of their mounting points on the legs, while the Body-IMU propagates the state of the main body. Simultaneously, the Leg-IMUs detect foot contact, i.e., zero velocity, to update their respective states. Additionally, relative position constraints between the Body-IMU and Leg-IMUs (estimated with joint encoders) are incorporated to mitigate error accumulation across all IMUs. We use the estimated IMU sensor errors to correct the IMU readings.
@@ III-C Observation Model
III-C1 Zero Velocity Update
@@
where indicates the noise. Consequently, we can build the zero velocity measurement model as
@@
This zero-velocity update is applied to the Leg-IMUs only when foot contact is detected. However, if all IMUs, including the Body-IMU, report zero velocity, the robot is considered to be in a static state, and the zero-velocity update is performed for all IMUs.
@@ III-C2 Relative Position Constraints
In our proposed DogLegs system, all the Body-IMU and Leg-IMUs perform IMU propagartion to predict their respective states. However, rather than moving independently, these IMUs are physically constrained within the robot’s kinematic chain. Their relative positions can be computed in real time using joint encoder readings and known robot parameters. This information helps to constrain the error drift of individual IMUs, improving overall state estimation accuracy.
Since the multiple IMUs are not aligned together, each has its own independent initial reference frame in which the IMU pose is expressed. As shown in Fig. 1, the world-frame is defined with the body-frame at the initial position. Therefore, to determine the Leg-IMU pose in the world-frame, we must compute the transformation between the initial Leg-IMU frame and the initial Body-IMU frame.
Taking one Leg-IMU as an example, its transformation w.r.t. the world-frame is represented by the rotation matrix and the translation vector . The roll and pitch components of can be determined by IMU data during the initial stationary period, while the translation can be obtained from leg kinematic model during the same period. For the yaw component of , we assume it to be zero, given that the IMUs are carefully installed and aligned during setup. Therefore, given this transformation and considering the position of the foot contact point in the Leg-IMU frame , the Leg-IMU can estimate the position of the foot in the world-frame, denoted as .
@@
where is the vector containing all joint angle measurements from the joint encoders of the leg, is the kinematics function of the leg, indicates the true position of the foot in the body-frame, and indicates the uncertainty of the relative position measurement, which is determined by the noise of the joint encoders. Consequently, we can build the relative position constraint model as
Whenever a Leg-IMU detects foot contact, we perform both zero velocity updates and relative position updates between that Leg-IMU and Body-IMU.
@@ IV-A Experimental Setup
We conducted experiments using a Unitree Go2 quadruped, which was equipped with a Body-IMU, four Leg-IMUs, a Global Navigation Satellite System (GNSS)/IMU integrated navigation system, and a raspberry pi for data collection, as shown in Fig. 3. The Body-IMU and Leg-IMUs were of the same type, containing a low-cost ICM20602 inertial sensor222https://invensense.tdk.com/download-pdf/icm-20602-datasheet/ (approximately 3 US dollars), a rechargeable battery, and a Bluetooth module for data transmission. The Leg-IMUs were mounted on the lower legs of the robot and close to the feet, while the Body-IMU was mounted on the robot body.
Sequence | Method | RPE tra. | RPE rot. | APE tra. | APE rot. |
Asphalt road | Foot-INS | ||||
LegOdom | |||||
DogLegs | |||||
Parking lots | Foot-INS | ||||
LegOdom | |||||
DogLegs | |||||
Construction site | Foot-INS | ||||
LegOdom | |||||
DogLegs | |||||
Offroad | Foot-INS | ||||
LegOdom | |||||
@@ IV-B State Estimation Performance
Because the Leg-IMUs have different roll and pitch angles to the main body, we assign the roll and pitch angles estimated by DogLegs to Foot-INS for a fair comparison. In addition, we report the mean value of the four individual systems on the four legs for Foot-INS. Table I lists the relative pose error (RPE) and absolute pose error (APE) calculated using evo [30] in four sequences. All the estimated trajectories were aligned with the ground truth using the transformation computed by evo [30] before metric evaluation.
Fig. 5 compares the estimated trajectories of LegOdom and our proposed DogLegs against the ground truth across the four sequences. To illustrate the actual absolute error drift, in this figure, we didn’t compute the transformation to align the whole estimated trajectories with the ground truth as done in evo [30]. As shown in the asphalt road sequence (a) and offroad sequence (d), DogLegs demonstrates significantly smaller error drift compared to LegOdom.
@@
Sequence | Method | Roll () | Pitch () | ||
RMSE | MAX | RMSE | MAX | ||
Asphalt road | LegOdom | ||||
DogLegs | |||||
Parking lots | LegOdom | ||||
DogLegs | |||||
Construction site | LegOdom | ||||
DogLegs | |||||
Offroad | LegOdom | ||||
DogLegs |

### v2

@@ Abstract
Robust and accurate proprioceptive state estimation of the main body is crucial for legged robots to execute tasks in extreme environments where exteroceptive sensors, such as LiDARs and cameras, may become unreliable. In this paper, we propose DogLegs, a state estimation system for legged robots that fuses the measurements from a body-mounted inertial measurement unit (Body-IMU), joint encoders, and multiple leg-mounted IMUs (Leg-IMU) using an extended Kalman filter (EKF). The filter system contains the error states of all IMU frames. The Leg-IMUs are used to detect foot contact, thereby providing zero-velocity measurements to update the state of the Leg-IMU frames. Additionally, we compute the relative position constraints between the Body-IMU and Leg-IMUs by the leg kinematics and use them to update the main body state and reduce the error drift of the individual IMU frames. Field experimental results have shown that our proposed DogLegs system achieves better state estimation accuracy compared to the traditional leg odometry method (using only Body-IMU and joint encoders) across various terrains. We make our datasets publicly available to benefit the research community (https://github.com/YibinWu/leg-odometry).
@@ I Introduction
Current approaches to this problem primarily use the robot body-mounted IMU (Body-IMU), joint encoders and foot force sensors [1, 8, 4, 9]. In these systems, the Body-IMU initially propagates the robot’s main body state, which is then refined by integrating leg kinematics estimated from joint encoders upon detecting foot contact using the foot force sensor. The most widely used approach for foot contact detection relies on empirically thresholding force sensor readings, but this method is susceptible to errors due to slippage and sensor degradation over time. In addition, the foot positions during swing phase cannot be updated [1, 2] and need to be relocated when the foot regains contact with the ground, introducing further error into the robot state estimation.
@@
The main contribution of this paper is the development of DogLegs, a framework that fuses the full state estimation information from all the individual IMU systems, including the Body-IMU and multiple Leg-IMUs, through an EKF. Specifically, each IMU system independently performs state propagation while a comprehensive error-state vector is maintained to account for errors across all IMU frames. The Leg-IMUs are used for foot contact detection, providing zero-velocity measurements to update the state of the Leg-IMU frames. Additionally, we use leg kinematics to compute the relative position constraints between the Body-IMU and Leg-IMUs. This measurement helps update the main body state, reduce error drift in individual IMU frames, and estimate and compensate for IMU bias. Fig. 1 illustrates the installation of multiple Leg-IMUs and the coordinate systems involved within our proposed DogLegs framework, using a quadruped robot as an example.
@@ II-A Proprioceptive Legged Robot State Estimation
An early approach to proprioceptive state estimation for legged robots was proposed by Bloesch et al. [1], who employed an EKF to fuse outputs from a Body-IMU, joint encoders, and foot force sensors. This method augmented the foot contact position into the state and updated it with the main body state using the leg kinematics when the foot contacts ground. This approach has become the basis for many subsequent state estimation frameworks. Later, Bloesch et al. [8] extended this work by incorporating the zero-velocity measurement and replacing the EKF with an unscented Kalman filter to improve accuracy and robustness.
@@ III Our Approach
Fig. 2 illustrates the algorithmic structure of our proposed DogLegs system. Our method is implemented using an error-state EKF to fuse the measurements from joint encoders and all the IMUs. The Leg-IMUs propagate the states of their mounting points on the legs, while the Body-IMU propagates the state of the main body. Simultaneously, the Leg-IMUs detect foot contact, i.e., zero-velocity, to update their respective states. Additionally, relative position constraints between the Body-IMU and Leg-IMUs (estimated with joint encoders) are incorporated to mitigate error accumulation across all IMUs. We use the estimated IMU sensor errors to correct the IMU readings.
@@ III-C Observation Model
III-C1 Zero-Velocity Update
@@
where indicates the noise. Consequently, we can build the zero-velocity measurement model as
@@
This zero-velocity update is applied to the Leg-IMUs only when foot contact is detected. However, if all IMUs, including the Body-IMU, report zero-velocity, the robot is considered to be in a static state, and the zero-velocity update is performed for all IMUs.
@@ III-C2 Relative Position Constraints
In our proposed DogLegs system, all the Body-IMU and Leg-IMUs perform the strapdown INS propagartion to predict their respective states. However, rather than moving independently, these IMUs are physically constrained within the robot’s kinematic chain [30]. Their relative positions can be computed in real time using joint encoder readings and known robot parameters. This information helps to constrain the error drift of individual IMUs, improving overall state estimation accuracy.
Since the multiple IMUs are not aligned together, each has its own independent initial reference frame in which the IMU pose is expressed. As shown in Fig. 1, the world-frame is defined with the body-frame at the initial position. Therefore, to determine the Leg-IMU’s pose in the world-frame, we must compute the transformation between the initial Leg-IMU frame and the initial Body-IMU frame, namely, world-frame.
Taking one Leg-IMU as an example, its transformation w.r.t. the world-frame is represented by the rotation matrix and the translation vector . The initial roll and pitch components of can be determined by IMU data during the initial stationary period of the robot, while the initial translation can be obtained from leg kinematic model during the same period. For the initial value of the yaw component of , we assume it to be zero, given that the IMUs are carefully installed and aligned during setup. Therefore, given this transformation and considering the position of the foot contact point in the Leg-IMU frame , the Leg-IMU can estimate the position of the foot in the world-frame, denoted as .
@@
where is the vector containing all joint angle measurements from the joint encoders of the leg, is the kinematics function of the leg [30], indicates the true position of the foot in the body-frame, and indicates the uncertainty of the relative position measurement, which is determined by the noise of the joint encoders and the uncertainty of the leg parameters. Consequently, we can build the relative position constraint model as
Whenever a Leg-IMU detects foot contact, we perform both zero-velocity updates and relative position updates between that Leg-IMU and Body-IMU.
@@ IV-A Experimental Setup
We conducted experiments using a Unitree Go2 quadruped, which was equipped with a Body-IMU, four Leg-IMUs, a Global Navigation Satellite System (GNSS)/IMU integrated navigation system, and a raspberry pi for data collection, as shown in Fig. 3. The Body-IMU and Leg-IMUs were of the same type, containing a low-cost ICM20602 inertial sensor222https://invensense.tdk.com/download-pdf/icm-20602-datasheet/ (approximately 3 US dollars), a rechargeable battery, and a Bluetooth module for data transmission. The Leg-IMUs were mounted on the lower legs of the robot and close to the feet, while the Body-IMU was mounted on the robot body. The extrinsic transformation between the Body-IMU frame and body-frame was measured in advance.
| Sequence | Method | RPE tra. | RPE rot. | APE tra. | APE rot. | |
Foot-INS | |||||
| LegOdom | ||||||
| DogLegs | ||||||
Foot-INS | |||||
| LegOdom | ||||||
| DogLegs | ||||||
Foot-INS | |||||
| LegOdom | ||||||
| DogLegs | ||||||
Foot-INS | |||||
| LegOdom | ||||||
@@ IV-B State Estimation Performance
We do not include a comparison with Yang et al. [2] because it drifted quickly and failed to produce meaningful results in our experiments. To ensure a fair comparison, we assign the roll and pitch angles estimated by DogLegs to Foot-INS, as the Leg-IMUs exhibit different roll and pitch orientations relative to the main body. In addition, we report the mean value of the four individual systems on the four legs for Foot-INS. Table I lists the relative pose error (RPE) and absolute pose error (APE) calculated using evo [31] in four sequences. All the estimated trajectories were aligned with the ground truth using the transformation computed by evo [31] before metric evaluation.
Roll () | Pitch () | ||||
| RMSE | MAX | RMSE | MAX | ||||
LegOdom | ||||||
| DogLegs | |||||||
LegOdom | ||||||
| DogLegs | |||||||
LegOdom | ||||||
| DogLegs | |||||||
LegOdom | ||||||
| DogLegs | |||||||
Fig. 5 compares the estimated trajectories of LegOdom and our proposed DogLegs against the ground truth across the four sequences. To illustrate the actual absolute error drift, in this figure, we didn’t compute the transformation to align the whole estimated trajectories with the ground truth as done in evo [31].
As shown in Table I, DogLegs achieves overall superior performance across all four sequences. However, in the parking lots sequence, it shows a slightly larger translation error than LegOdom, which is likely due to the stochastic error of the inertial sensors. Furthermore, the alignment of the estimated trajectory with the ground truth may partially compensate for actual drift, thereby obscuring the true performance advantage of the proposed DogLegs. This is illustrated in Fig. 5 (b), where the trajectory of DogLegs more closely matches the ground truth compared to LegOdom.
In addition, Fig. 5 (a) and (d) clearly show that DogLegs exhibits significantly smaller drift compared to LegOdom. Particularly, in the offroad sequence shown in Fig. 5 (d), LegOdom exhibits a large heading error at the beginning of the trajectory, leading to a large drift in the estimated trajectory. In contrast, DogLegs maintains a more stable heading estimation, resulting in a more accurate trajectory.
@@ IV-D Foot Contact Detection
IV-E Computation Efficiency
In this section, we analyze the computational efficiency of our proposed DogLegs. All experiments were conducted on a desktop computer equipped with an Intel i7-10700 CPU running at and of RAM, using a single computation thread. The average computation time per IMU measurement of DogLegs across all sequences is summarized in Table III. The prediction time includes the strapdown INS propagation for all IMUs as well as the error-state propagation in the EKF. The update time accounts for the processing of all zero-velocity updates and relative position constraints. Note that multiple foot contact events may occur simultaneously. The total time represents the overall time required to process all the IMU measurements and joint encoder received coming at the same time. All values are reported in milliseconds (ms) in Table III.
As shown in Table III, the average computation time of DogLegs is approximately per measurement, which is significantly lower than the IMU measurement interval of at . This indicates that DogLegs can run in real-time on a standard desktop computer. The prediction step takes about , while the update step takes about . The longer duration of the update step is due to the additional computations involved in processing zero-velocity updates and relative position constraints triggered by all foot contact events. Notably, the average total processing time is shorter than the average update time because the update step is not executed at every time step, namely, it is only performed when a foot contact is detected.
| Prediction | Update | Total |